A steel pipe 2ft in diameter carries water at about 15ft/s. If the pipe has a roughness of 0.0003ft, could the capacity be increased by inserting a smooth plastic liner that reduces the inside diameter to 1.9ft ? Calculate the change in pressure drop for the same flow and the change in capacity for fixed pressure drop. (In fluid mechanics)
1
Expert's answer
2020-04-03T11:48:38-0400

d=2ft

rate/pressure=15ft/s

thickness=0.003ft

If the thickness is increased to 1.9ft the diameter will be decreased hence reducing the capacity of the pipe per length.

hence the pressure,p becomes;

"p=area\/d"

area=15*2=30ft

p=30/(2-1.9)

p=300ft/s